### v1

@@ 3.3 Incorrect Ground Truth and Ambiguity
Our choice of a frontier model as a judge is motivated by prior work suggesting that models are significantly stronger verifiers than generators (verifier_Liao_2025_CVPR; verifier_saad2025shrinking; verifier_venktesh2025trustverifysurveyverification; verifier-guan-etal-2024-language); we therefore expect the judge to accurately identify errors even in cases that are too challenging for contemporary models to solve autonomously. Given that we operate in a regime of abundant evaluation data across our 9 capabilities, we intentionally err on the side of caution. We adopt a stringent filtering policy, discarding any item flagged as (1) ambiguous, (2) incorrectly labeled, or (3) unsolvable due to insufficient resolution, ensuring that the resulting DatBench subset represents only the highest quality of evaluation data. The impact of this pipeline is most visible in the Spatial capability, which exhibits a 42.07% discard rate, primarily due to insufficient resolution in “in-the-wild” images. Similarly, complex expert-authored sets like ChartQA Pro (17.2% removed) and MMMU-Pro (24.3% removed) show significantly higher noise rates than standard benchmarks (c.f. Appendix D for per dataset / capability counts of filtered examples). While these high attrition rates reflect significant noise in frontier evaluations, we recognize that a judge might occasionally misinterpret specialized, valid reasoning as a data defect. To maintain evaluative headroom, we retain only the subset of these examples that the judge explicitly validates as correct and unambiguous. Our aggregate data surplus allows us to prioritize this high-fidelity subset, accepting the risk that a conservative filtering policy may sacrifice some valid frontier samples to ensure the remaining benchmark remains strictly noise-free.
@@ 3.4 High Discrimination with Limited Compute
As VLMs grow in sophistication and expand their set of capabilities, comprehensive evaluation imposes a prohibitive computational burden. This is exacerbated by the emergence of “thinking” models; for instance, qwen3vl utilize inference-time compute scaling, often generating chains-of-thought exceeding 32K tokens. Consequently, evaluating a single capability like OCR (often containing >100>100K examples) can require generating over 3 billion tokens, an untenable cost for iterative research.
Selecting a representative subset of examples is a natural approach to reducing evaluation costs. The intuitive heuristic for such a selection is to preserve the model ranking induced by the full dataset, typically quantified using rank correlation measures such as Spearman’s $\rho$ or Kendall’s $\tau$ (spearman1904; voorhees2001; buckley2004). While rank preservation is a necessary condition for a representative subset, it is theoretically insufficient: rank correlation is agnostic to which specific samples are retained. In practice, even random subsets can preserve global model rankings by retaining items that separate coarse capability tiers (e.g., small versus large models), while failing to retain the high-discrimination examples needed to distinguish models along the Pareto frontier. More broadly, methods that optimize solely for rank preservation face a fundamental limitation, rank correlation saturates rapidly and is often achieved by subsets whose individual samples are weakly or inconsistently informative about underlying capabilities (sakai2007; voorhees2001). In such regimes, apparent ranking stability may be driven by spurious correlations or superficial artifacts rather than genuine reasoning ability.
Instead, we turn to Item Response Theory (IRT) for inspiration, originally formalized by lord1952theory. IRT posits that items differ not just in difficulty, but in item discrimination, a parameter that determines how sharply an item distinguishes between subjects of varying ability levels (baker2001basics). However, directly applying standard IRT methodologies (fluid2025) to VLM evaluation is often infeasible due to the limited number of diverse observations available per sample in the current research landscape (bean2025scales++; liao2025unifiedframeworkdataefficientevaluation). Effectively fitting IRT models typically requires stable evaluations from hundreds of diverse state-of-the-art models; without this scale, IRT models become highly sensitive to hyperparameters and are notoriously difficult to fit stably.
@@ 5 Diagnosing VLM Pathologies with DatBench
Comparing Thinking models to standard counterparts reveals that extra test-time compute is a double-edged sword (Figure 11(a)). To quantify this, we define the Thinking relative advantage as the percentage gain in accuracy of the thinking model over its instruct counterpart, normalized by the instruct baseline: $(Acc_{thinking}-Acc_{instruct})/Acc_{instruct}\times 100$. Scaling helps Math ($\approx+36.8\%$) and Charts ($\approx+10.8\%$) but causes massive regressions in OCR ($\approx-53.5\%$) and Document Understanding ($\approx-47.8\%$). This regression is also extremely computationally wasteful: while correct thinking answers use $\approx 425$ tokens, incorrect attempts balloon to $\approx 1196.9$ tokens, a $\approx 14\times$ increase over non-thinking models (Figure 11(b)). We observed that this is due to models entering unproductive thinking loops on perceptual tasks they cannot solve. Prior work has observed a similar overthinking penalty for language models (overthink_hochlehnert2025soberlookprogresslanguage; overthink_su2025underthinkingoverthinkingempiricalstudy; overthink_wang2025thoughtsplaceunderthinkingo1like; overthink_wu2025lessunderstandingchainofthoughtlength).

### v2

@@ 3.3 Incorrect Ground Truth and Ambiguity
Our choice of a frontier model as a judge is motivated by prior work suggesting that models are significantly stronger verifiers than generators (Liao et al., 2025b; Saad-Falcon et al., 2025; Venktesh et al., 2025; Guan et al., 2024); we therefore expect the judge to accurately identify errors even in cases that are too challenging for contemporary models to solve autonomously. Given that we operate in a regime of abundant evaluation data across our 9 capabilities, we intentionally err on the side of caution. We adopt a stringent filtering policy, discarding any item flagged as (1) ambiguous, (2) incorrectly labeled, or (3) unsolvable due to insufficient resolution, ensuring that the resulting DatBench subset represents only the highest quality of evaluation data. The impact of this pipeline is most visible in the Spatial capability, which exhibits a 42.07% discard rate, primarily due to insufficient resolution in “in-the-wild” images. Similarly, complex expert-authored sets like ChartQA Pro (17.2% removed) and MMMU-Pro (24.3% removed) show significantly higher noise rates than standard benchmarks (c.f. Appendix D for per dataset / capability counts of filtered examples). While these high attrition rates reflect significant noise in frontier evaluations, we recognize that a judge might occasionally misinterpret specialized, valid reasoning as a data defect. To maintain evaluative headroom, we retain only the subset of these examples that the judge explicitly validates as correct and unambiguous. Our aggregate data surplus allows us to prioritize this high-fidelity subset, accepting the risk that a conservative filtering policy may sacrifice some valid frontier samples to ensure the remaining benchmark remains strictly noise-free.
@@ 3.4 High Discrimination with Limited Compute
As VLMs grow in sophistication and expand their set of capabilities, comprehensive evaluation imposes a prohibitive computational burden. This is exacerbated by the emergence of “thinking” models; for instance, Bai et al. (2025) utilize inference-time compute scaling, often generating chains-of-thought exceeding 32K tokens. Consequently, evaluating a single capability like OCR (often containing >100>100K examples) can require generating over 3 billion tokens, an untenable cost for iterative research.
Selecting a representative subset of examples is a natural approach to reducing evaluation costs. The intuitive heuristic for such a selection is to preserve the model ranking induced by the full dataset, typically quantified using rank correlation measures such as Spearman’s $\rho$ or Kendall’s $\tau$ (Spearman, 1904; Voorhees, 2001; Buckley and Voorhees, 2004). While rank preservation is a necessary condition for a representative subset, it is theoretically insufficient: rank correlation is agnostic to which specific samples are retained. In practice, even random subsets can preserve global model rankings by retaining items that separate coarse capability tiers (e.g., small versus large models), while failing to retain the high-discrimination examples needed to distinguish models along the Pareto frontier. More broadly, methods that optimize solely for rank preservation face a fundamental limitation, rank correlation saturates rapidly and is often achieved by subsets whose individual samples are weakly or inconsistently informative about underlying capabilities (Sakai, 2007; Voorhees, 2001). In such regimes, apparent ranking stability may be driven by spurious correlations or superficial artifacts rather than genuine reasoning ability.
Instead, we turn to Item Response Theory (IRT) for inspiration, originally formalized by Lord (1952). IRT posits that items differ not just in difficulty, but in item discrimination, a parameter that determines how sharply an item distinguishes between subjects of varying ability levels (Baker, 2001). However, directly applying standard IRT methodologies (Schick and others, 2025) to VLM evaluation is often infeasible due to the limited number of diverse observations available per sample in the current research landscape (Bean et al., 2025; Liao et al., 2025a). Effectively fitting IRT models typically requires stable evaluations from hundreds of diverse state-of-the-art models; without this scale, IRT models become highly sensitive to hyperparameters and are notoriously difficult to fit stably.
@@ 5 Diagnosing VLM Pathologies with DatBench
Comparing Thinking models to standard counterparts reveals that extra test-time compute is a double-edged sword (Figure 11(a)). To quantify this, we define the Thinking relative advantage as the percentage gain in accuracy of the thinking model over its instruct counterpart, normalized by the instruct baseline: $(Acc_{thinking}-Acc_{instruct})/Acc_{instruct}\times 100$. Scaling helps Math ($\approx+36.8\%$) and Charts ($\approx+10.8\%$) but causes massive regressions in OCR ($\approx-53.5\%$) and Document Understanding ($\approx-47.8\%$). This regression is also extremely computationally wasteful: while correct thinking answers use $\approx 425$ tokens, incorrect attempts balloon to $\approx 1196.9$ tokens, a $\approx 14\times$ increase over non-thinking models (Figure 11(b)). We observed that this is due to models entering unproductive thinking loops on perceptual tasks they cannot solve. Prior work has observed a similar overthinking penalty for language models (Hochlehnert et al., 2025; Su et al., 2025; Wang et al., 2025; Wu et al., 2025).